b) Given the functions f(x) is 3x^2-5 and g(x) is 4+2x, find the following
(i) (fog) (2) (ii) g(^-1)

Answer by Boreal(15235)

f o g=3(4+2x)^2-5
can put 2 in right here for x
3(4+4)^2-5
3(64)-5=187.
--------------
g^(-1):
y=4+2x
x=2+4y, solve for y
4y=x-2
y=(1/4)(x-2)=g^(-1)